OpenStudy (anonymous):

Diane flipped a coin 600 times and it landed heads up 240 times. What is the relative frequency of the coin landing heads up based on the results of this experiment? 4% 6% 40% 60%

OpenStudy (anonymous):

40%

OpenStudy (anonymous):

240/600 -> 24/60 / 6 = 4/10 x 10 = 40/100